OpenStudy (anonymous):

HI CAN YOU answer this.. f(x)=3x^2-2x+5 by answering completing the squares and getting the h and k!!!! please needs answers now

OpenStudy (anonymous):

???????????

OpenStudy (wolf1728):

3x^2-2x+5 = 0 Move non-x term to the right 3x^2-2x = -5 Divide equation by coefficient of x^2 x^2 -(2x/3) = -5/3 Get coefficient of x (-2/3) divide it by 2 then square it. -2/3 divided by 2 = -1/3 squaring -1/3 equals 1/9 then add to both sides of equation x^2 -(2x/3) +1/9 = -5/3 +1/9 Take square root of both sides (x + 1/3) = Sqr Root (-15/9 + 1/9) (x + 1/3) = Sq Root (-14/9) ****************************************************************** Don't know about negative square roots but the answers are: 0.33333 plus 1.2472 i and 0.33333 minus 1.2472 i

OpenStudy (anonymous):

thnks.....

OpenStudy (wolf1728):

u r welcome - hope it helps

OpenStudy (anonymous):

wait a minute

OpenStudy (anonymous):

when you divide 2 to -2/3 it seems like -2/3/3 its equal to -2/6 then you will sqare it it will be 4/36

OpenStudy (anonymous):

please i need answer